THE ELECTRONIC SPECTRUM AND MOLECULAR STRUCTURE OF THE ARSENYL (H_{2}As=O) FREE RADICAL

SHENG-GUI HE, Institute of Chemistry, Chinese Academy of Sciences, Zhongguancun, Beijing 100080, P.R.C.; \underline{FUMIE X. SUNAHORI}, JIE YANG, AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The H_{2}As=O radical has been identified for the first time by laser-induced fluorescence (LIF) and single vibronic level (SVL) emission techniques. The radical was generated by a pulsed electric discharge in a mixture of AsH_{3} and CO_{2} and high-pressure argon and detected by observation of the B̃^{2}A′ − X̃^{2}A′ electronic transition in the 510-410 nm region. Moderate resolution LIF and SVL emission spectra of H_{2}AsO, D_{2}AsO, and HDAsO have been recorded and analysis shows unequivocally that the spectrum is due to the arsenyl radical. High-resolution spectra of the 0_{0}^{0} bands of H_{2}AsO and D_{2}AsO, which consist of strong a-type and weaker c-type transitions, revealed spin-splittings and small, but significant arsenic hyperfine splittings due to a Fermi contact interaction in the ground state. The effective molecular structures of H_{2}AsO in the ground and excited states have been determined from the rotational constants and will be discussed in the context of the analogous nitroxyl (X_{2}N=O) and phosphonyl (X_{2}P=O) radicals.